Charges of 7.2nC and 6.7nC are 32 cm apart. Find the equilibrium position for a -3.0nC charge?????


1
Expert's answer
2021-03-14T19:20:51-0400

"F_1=k\frac{q_1q}{x^2}"


"F_2=k\frac{q_2q}{(0.32-x)^2}"


"F_1=F_2\to k\frac{q_1q}{x^2}=k\frac{q_2q}{(0.32-x)^2}\to\frac{q_1}{x^2}=\frac{q_2}{(0.32-x)^2}\to"


"q_1(0.32-x)^2=q_2x^2\to(q_1-q_2)x^2-0.64q_1x+0.1024q_1=0"


"0.5x^2-4.608x+0.73728=0\to x\approx0.163\ (m)=16.3\ (cm)" . Answer